# Interplay between local and global order controls bifurcations in nonreciprocal systems without a conservation law

### v2

@@ Abstract
Conclusion.—We explained the collective dynamics of the nonreciprocal Ising system on the level of both, local and global order. A critical threshold magnitude of local order within the respective lattices was found to control the emergence of coherent oscillations of the global order parameter. Upon increasing interactions, ghost states emerge and the residence time in either of them eventually diverges, giving rise to a non-equilibrium steady state via a saddle-node-infinite-period bifurcation. Strikingly, during coherent oscillations of global order, a high degree of local order is preserved in nearest-neighbor-interacting systems. This implies non-trivial spatio-temporal correlations between spins, confirmed by a spectral-density maximum at large-scale low-frequency modes. In stark contrast, on the mean-field (all-to-all) lattice there is no distinction between different spatial modes, annihilating any notion of spatial structure. Thus, accounting for nearest-neighbor correlations is essential for a correct understanding of dynamics of nonreciprocal matter with short, or more generally finite, range of interactions.
Our work provides a comprehensive microscopic understanding of dynamic collective phenomena in nonreciprocal matter without conservation laws relevant, e.g. for enzyme clustering [67, 68], where the nonreciprocal coupling relates to an attractive $K>0$ or repulsive $K<0$ interaction between two enzymes where the product of one is the substrate of the other enzyme. The magnetizations $m^{\mu}(t)$ in turn reflect time-dependent concentrations of product and substrates via $\phi^{\mu}(t)=(1+m^{\mu}(t))/2\in[0,1]$.
@@
What remains elusive are multiple ($>2$) coupled lattices, spatially heterogeneous/extended systems [69], as well as the thermodynamic cost of dynamical states and bifurcations [53, 70, 71, 72]. Moreover, considering the relevance of conservation laws [73, 2, 41], it will be essential to develop a theoretical framework for the nonreciprocal Ising model with Kawasaki dynamics. These will be addressed in future works.
@@
Acknowledgments.— Financial support from the German Research Foundation (DFG) through the Heisenberg Program Grants GO 2762/4-1 and GO 2762/5-1 (to AG) is gratefully acknowledged.
@@
Appendix A: Mean-field approximation.—A less accurate technique to obtain approximate evolution equations is the mean-field (MF) approximation [48], where one makes the rudimentary (uncontrolled) assumption
$\langle\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\rangle\approx\tanh{\langle\Delta E^{\mu}_{i}/2\rangle},$ (A1)
@@
$\tau\frac{dm^{a}(t)}{dt}+m^{a}(t)$ $=\tanh{(zJ_{a}m^{a}(t)+K_{a}m^{b}(t))},$
$\tau\frac{dm^{b}(t)}{dt}+m^{b}(t)$ $=\tanh{(zJ_{b}m^{b}(t)+K_{b}m^{a}(t))},$ (A2)
which are exact on the fully connected mean-field lattice, where the local order is trivial (i.e. there is no sense of “local”), $q^{\mu\nu}(t)=m^{\mu}(t)m^{\nu}(t)$, and therefore $\mathcal{C}^{\mu\nu}(t)=0$. A linear stability analysis around the trivial steady-state $m^{\mu}_{s}=0$ for $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$ leads to the linear stability equation $\tau d\delta\mathbf{m}(t)/dt=\mathbf{M}^{\rm MF}(J,K)\delta\mathbf{m}(t)$ with elements
$M^{\rm MF}_{aa}$ $=M^{\rm MF}_{bb}=zJ-1,$ (A3)
$M^{\rm MF}_{ab}$ $=-M^{\rm MF}_{ba}=K.$ (A4)
The eigenvalues of $\mathbf{M}^{\rm MF}$ are given by
$\lambda^{\rm MF}_{\pm}(J,K)=(zJ-1)\pm iK.$ (A5)
Hence, the Hopf bifurcation occurs at $J=1/z$ and $K\neq 0$, such that ${\rm Re}(\lambda^{\rm MF}_{\pm})=0$ and ${\rm Im}(\lambda^{\rm MF}_{\pm})\neq 0$. This corresponds to a straight vertical line in the $(J,K)$-plane, as shown in Fig. A1a, which is unphysical for short range interactions. Notably, in the MF approximation we do not observe a critical value for local order, which is present in the more accurate BG approximation.
Figure A1: Phase diagram for global order $m^{\mu}_{s}$ obtained with the mean-field (MF) approximation (a) and the monomer approximation (b) for the perfect nonreciprocal setting with $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$. In the MF approximation the Hopf bifurcation (blue line) is set by $J=1/z$. In the monomer approximation, the Hopf bifurcations (blue line) is set by a critical local order $q_{s}=1/z$, which is more similar to the Bethe-Guggenheim approximation where $q_{s}=1/(z-1)$ (see Fig. 2d).
@@
MF (A2) $=m^{\mu}(t)m^{\nu}(t)$ $J{=}1/z$
monomer (5)+(B2) (5)+(B2); slaved by $m^{\mu}(t)$ $q_{s}{=}1/z$
BG (5)+(6) (5)+(6); not slaved $q_{s}{=}1/(z{-}1)$
Appendix B: Monomer approximation.—Another approximation technique we developed in this work is what we call the “monomer approximation”. It is more accurate than the MF yet less accurate than the BG approximation.
The conceptual difference between the MF on the one hand, and the monomer and BG approximations on the other hand lies in the treatment of the average $\langle\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\rangle$. Whereas the MF approximation simply moves the average to the argument as shown in Eq. (A1), the monomer and BG approximations use the fact that the value of $\Delta E^{\mu}_{i}/2$ lies in an enumerable set given by $U^{\mu}_{l,n}\equiv[2l-z]J_{\mu}+[2n-1]K_{\mu}$ with $l\in\{0,..,z\}$ and $n\in\{0,1\}$. This allows for an explicit summation
$\langle\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\rangle=\sum_{l=0}^{z}\sum_{n=0}^{1}\mathcal{P}^{\mu}_{l,n}(t)\tanh{(U^{\mu}_{l,n})},$ (B1)

### v3

@@ VI Concluding remarks
We explained the collective dynamics of the nonreciprocal Ising system on the level of both, local and global order beyond the mean-field approximation. A critical threshold magnitude of local order within the respective lattices was found to control the emergence of coherent oscillations of the global order parameter. Upon increasing interactions, ghost states emerge and the residence time in either of them eventually diverges, giving rise to a dynamically trapped terminal state via a saddle-node-infinite-period bifurcation. The terminal state depends on the initial condition; the dynamics in this regime is thus non-ergodic.
Strikingly, during coherent oscillations of global order, a high degree of local order is preserved (see Fig. 1d). This implies nontrivial spatiotemporal correlations between spins, confirmed by a spectral-density maximum at large-scale low-frequency modes. In stark contrast, on the mean-field (all-to-all) lattice there is no distinction between different spatial modes, annihilating any notion of spatial structure. Thus, accounting for nearest-neighbor correlations is essential for a correct understanding of dynamics of nonreciprocal matter with short, or more generally finite, range of interactions.
@@
Our work provides a comprehensive microscopic understanding of dynamic collective phenomena in nonreciprocal matter without conservation laws based on the nonreciprocal Ising model. What remains elusive are multiple ($>2$) coupled lattices, spatially heterogeneous/extended systems Blom et al. (2023), as well as the thermodynamic cost of dynamical states and bifurcations Suchanek et al. (2023); Nardini et al. (2017); Blom et al. (2024); Zhang and Garcia-Millan (2023). Moreover, considering the relevance of conservation laws Frohoff-Hülsmann and Thiele (2023); Brauns and Marchetti (2024); Saha et al. (2020); Greve and Thiele (2024), it will be essential to develop a theoretical framework for the nonreciprocal Ising model with Kawasaki dynamics. These will be addressed in future work.
VII Acknowledgments
@@
Financial support from the German Research Foundation (DFG) through the Heisenberg Program Grants GO 2762/4-1 and GO 2762/5-1 (to AG) is gratefully acknowledged.
Appendix A Derivation of Eqs. (8)-(10)
From the master equation (2), we directly obtain dynamical equations for the first two moments of single-spin values (see also Eqs. (28) and (29) in Glauber (1963))
$\tau\frac{{\rm d}\langle\sigma^{\mu}_{i}(t)\rangle}{{\rm d}t}+\langle\sigma^{\mu}_{i}(t)\rangle$ $=\langle\sigma^{\mu}_{i}(t)\tanh{(\Delta E^{\mu}_{i}/2)}\rangle,$ (A1)
$\tau\frac{{\rm d}\langle\sigma^{\mu}_{i}(t)\sigma^{\nu}_{j}(t)\rangle}{{\rm d}t}+2\langle\sigma^{\mu}_{i}(t)\sigma^{\nu}_{j}(t)\rangle$ $=\langle\sigma^{\mu}_{i}(t)\sigma^{\nu}_{j}(t)[\tanh{(\Delta E^{\mu}_{i}/2)}+\tanh{(\Delta E^{\nu}_{j}/2)}]\rangle,$ (A2)
where $\langle f(t)\rangle\equiv\sum_{\boldsymbol{\sigma}}P(\boldsymbol{\sigma};t)f(\boldsymbol{\sigma})$. Eqs. (A1)-(A2) are exact (but not yet closed) equations for the first two moments evolving under Glauber dynamics, and will serve as our starting point to derive equations for the global and local order parameters, which we derive in two steps: First, we sum Eqs. (A1)-(A2) over all spins and spin pairs. Upon doing this, the left-hand side of Eqs. (A1)-(A2) transforms into
$\frac{1}{N}\sum_{i=1}^{N}\left(\tau\frac{{\rm d}\langle\sigma^{\mu}_{i}(t)\rangle}{{\rm d}t}+\langle\sigma^{\mu}_{i}(t)\rangle\right)$ $=\tau\frac{{\rm d}m^{\mu}(t)}{{\rm d}t}+m^{\mu}(t),$ (A3)
$\frac{2}{zN}\sum_{i=1}^{N}\sum_{\langle i|j\rangle}\left(\tau\frac{{\rm d}\langle\sigma^{\mu}_{i}(t)\sigma^{\mu}_{j}(t)\rangle}{{\rm d}t}+2\langle\sigma^{\mu}_{i}(t)\sigma^{\mu}_{j}(t)\rangle\right)$ $=\tau\frac{{\rm d}q^{\mu\mu}(t)}{{\rm d}t}+2q^{\mu\mu}(t),$ (A4)
$\frac{1}{N}\sum_{i=1}^{N}\left(\tau\frac{{\rm d}\langle\sigma^{a}_{i}(t)\sigma^{b}_{i}(t)\rangle}{{\rm d}t}+2\langle\sigma^{a}_{i}(t)\sigma^{b}_{i}(t)\rangle\right)$ $=\tau\frac{{\rm d}q^{ab}(t)}{{\rm d}t}+2q^{ab}(t).$ (A5)
Second, we need to evaluate the right-hand side of Eqs. (A1)-(A2) after summation over all spins and spin pairs. To do this, we note that $\Delta E^{\mu}_{i}$ can take on a discrete (enumerable) set of values. Consider a spin with $l\in\{0,1,...,z\}$ neighboring up spins on the same lattice and $n\in\{0,1\}$ neighboring up spins on the opposing lattice. We want to compute the change in energy upon flipping this spin. Based on Eq. (1) we can parameterize this change in energy upon flipping the spin as
$\Delta E^{\mu}_{i}=2\sigma^{\mu}_{i}U^{\mu}_{l,n},$ (A6)
where $U^{\mu}_{l,n}$ is given by Eq. (11). Using this parameterization, we evaluate the right-hand side of Eqs. (A1)-(A2)
$\frac{1}{N}\sum_{i=1}^{N}\langle\sigma^{\mu}_{i}\tanh{(\Delta E^{\mu}_{i}/2)}\rangle$ $=\frac{1}{N}\sum_{i=1}^{N}\langle\tanh{(U^{\mu}_{l,n})}\rangle$ $=\sum_{l=0}^{z}\sum_{n=0}^{1}\mathcal{P}^{\mu}_{l,n}(t)\tanh{(U^{\mu}_{l,n})},$ (A7)
$\frac{2}{zN}\sum_{i=1}^{N}\sum_{\langle i|j\rangle}\langle\sigma^{\mu}_{i}\sigma^{\mu}_{j}\tanh{(\Delta E^{\mu}_{i}/2)}\rangle$ $=\frac{2}{zN}\sum_{i=1}^{N}\sum_{\langle i|j\rangle}\langle\sigma^{\mu}_{j}\tanh{(U^{\mu}_{l,n})}\rangle$ $=\frac{2}{z}\sum_{l=0}^{z}\sum_{n=0}^{1}(2l-z)\mathcal{P}^{\mu}_{l,n}(t)\tanh{(U^{\mu}_{l,n})},$ (A8)
$\frac{1}{N}\sum_{i=1}^{N}\langle\sigma^{a}_{i}\sigma^{b}_{i}\tanh{(\Delta E^{a}_{i}/2)}\rangle$ $=\frac{1}{N}\sum_{i=1}^{N}\langle\sigma^{b}_{i}\tanh{(U^{a}_{l,n})}\rangle$ $=\sum_{l=0}^{z}\sum_{n=0}^{1}(2n-1)\mathcal{P}^{\mu}_{l,n}(t)\tanh{(U^{a}_{l,n})}.$ (A9)
For the first equality in Eqs. (A7)-(A9) we used $\tanh{(\Delta E^{\mu}_{i}/2)}=\tanh{(\sigma^{\mu}_{i}U^{\mu}_{l,n})}=\sigma^{\mu}_{i}\tanh{(U^{\mu}_{l,n})}$, together with $(\sigma^{\mu}_{i})^{2}=1$. For the second equality, we used that terms such as $\langle\tanh{(U^{\mu}_{l,n})}\rangle$ represent a weighted sum over all possible combinations of the possible values that $\tanh{(U^{\mu}_{l,n})}$ can attain. The weights are given by the time-dependent probability $\mathcal{P}^{\mu}_{l,n}(t)$ to find an up or down spin with a specific local environment. By definition, this probability is normalized $\sum_{l,n}\mathcal{P}^{\mu}_{l,n}(t)=1$. Combining Eqs. (A3)-(A5) and (A7)-(A9) we obtain Eqs. (8)-(10).
Appendix B Derivation of Eq. (14)
Here, we derive Eq. (14) based on the Bethe-Guggenheim (BG) approximation. We focus on the probability of picking a spin on the $a$ lattice with a given specific local environment. The same reasoning will also apply for picking a spin on the $b$ lattice. Recall that $\mathcal{P}^{a\pm}_{l,n}(t)$ is the probability at time $t$ to find an up ($+$) or down (--) spin with $l$ neighboring up spins on the $a$ lattice and $n$ neighboring up spins on the $b$ lattice. On the BG level, we assume ideal mixing of nearest-neighbor spin pairs, resulting in the following expressions
$\mathcal{P}^{a+}_{l,n}$ $=\underbrace{[N^{a}_{+}/N]}_{\begin{subarray}{c}\rm probability\\ \begin{subarray}{c}\rm for\ up\ spin\\ {\rm on\ the}\ a\ {\rm lattice}\end{subarray}\end{subarray}}\times\underbrace{\left[\binom{N^{aa}_{++}}{l}\binom{N^{aa}_{+-}/2}{z-l}/\binom{N^{aa}_{++}+N^{aa}_{+-}/2}{z}\right]}_{\begin{subarray}{c}\rm probability\ for\\ \begin{subarray}{c}l\ {\rm\ neighboring\ up}\\ {\rm spins\ on\ the}\ a\ {\rm lattice}\end{subarray}\end{subarray}}\times\underbrace{\left[\binom{N^{ab}_{++}}{n}\binom{N^{ab}_{+-}}{1-n}/\binom{N^{ab}_{++}+N^{ab}_{+-}}{1}\right]}_{\begin{subarray}{c}\rm probability\ for\\ \begin{subarray}{c}n\ {\rm\ neighboring\ up}\\ {\rm spins\ on\ the}\ b\ {\rm lattice}\end{subarray}\end{subarray}},$ (B1)
$\mathcal{P}^{a-}_{l,n}$ $=\underbrace{[N^{a}_{-}/N]}_{\begin{subarray}{c}\rm probability\\ \begin{subarray}{c}\rm for\ down\ spin\\ {\rm on\ the}\ a\ {\rm lattice}\end{subarray}\end{subarray}}\times\underbrace{\left[\binom{N^{aa}_{-+}/2}{l}\binom{N^{aa}_{--}}{z-l}/\binom{N^{aa}_{-+}/2+N^{aa}_{--}}{z}\right]}_{\begin{subarray}{c}\rm probability\ for\\ \begin{subarray}{c}l\ {\rm\ neighboring\ up}\\ {\rm spins\ on\ the}\ a\ {\rm lattice}\end{subarray}\end{subarray}}\times\underbrace{\left[\binom{N^{ab}_{-+}}{n}\binom{N^{ab}_{--}}{1-n}/\binom{N^{ab}_{-+}+N^{ab}_{--}}{1}\right]}_{\begin{subarray}{c}\rm probability\ for\\ \begin{subarray}{c}n\ {\rm\ neighboring\ up}\\ {\rm spins\ on\ the}\ b\ {\rm lattice}\end{subarray}\end{subarray}},$ (B2)
$2N^{\mu\mu}_{\pm\pm}+N^{\mu\mu}_{+-}$ $=zN^{\mu}_{\pm},$ (B3)
$N^{ab}_{\pm\pm}+N^{ab}_{\pm\mp}$ $=N^{a}_{\pm},$ (B4)
in combination with
$N^{\mu}_{\pm}$ $=N(1\pm m^{\mu})/2.$ (B5)
Furthermore, we use the definition of local order given by Eqs. (5)-(6) to write
$q^{\mu\mu}$ $=2(N^{\mu\mu}_{++}+N^{\mu\mu}_{--}-N^{\mu\mu}_{+-})/zN$
$=1-4N^{\mu\mu}_{+-}/zN,$ (B6)
$q^{ab}$ $=(N^{ab}_{++}+N^{ab}_{--}-N^{ab}_{+-}-N^{ab}_{-+})/N$
$=1-2N^{ab}_{+-}/N-2N^{ab}_{-+}/N$
$=1+m^{a}-m^{b}-4N^{ab}_{+-}/N,$ (B7)
where in the last line we used the relation
$m^{b}-m^{a}=2(N^{ab}_{-+}-N^{ab}_{+-})/N.$ (B8)
Using the relations (B3)-(B7) we obtain the following expression for the spin pairs within the same lattice
$N^{\mu\mu}_{\pm\pm}$ $=(z/8)N(1\pm 2m^{\mu}+q^{\mu\mu}),$ (B9)
$N^{\mu\mu}_{\pm\mp}$ $=(z/4)N(1-q^{\mu\mu}),$ (B10)
and for the spin pairs between the two opposing lattices
$N^{ab}_{\pm\pm}$ $=(1/4)N(1\pm m^{a}\pm m^{b}+q^{ab}),$ (B11)
$N^{ab}_{\pm\mp}$ $=(1/4)N(1\pm m^{a}\mp m^{b}-q^{ab}).$ (B12)
Inserting Eqs. (B9)-(B12) into Eqs. (B1)-(B2) and taking the thermodynamic limit $N\rightarrow\infty$ while keeping $m^{\mu}(t)$, $q^{\mu\mu}(t)$, and $q^{ab}(t)$ fixed, we obtain Eq. (14).
Appendix C Steady-state local order
The trivial steady state is given by the disordered state with $m^{\mu}_{s}=0$ and $q^{ab}_{s}=0$. To solve for the steady state of the local order, denoted as $q_{s}(J,K)$, we need to solve
$q_{s}(J,K)=\frac{1}{z}\sum_{l=0}^{z}\sum_{n=0}^{1}(2l-z)(\overline{\mathcal{P}}^{+}_{l}+\overline{\mathcal{P}}^{-}_{l})\tanh{(U^{a}_{l,n})},$ (C1)
where $\overline{\mathcal{P}}^{\pm}_{l}(q_{s})$ are the probabilities (14) evaluated at steady-state values given by Eq. (19). Equation (C1) can be solved for specific integer values of $z$. For example, for $z=2$ we obtain
$q_{s}(J,K)|_{z=2}=\frac{2-\sqrt{4-\left[\sum_{n=\pm}\tanh{(2J+nK)}\right]^{2}}}{\sum_{n=\pm}\tanh{(2J+nK)}}.$
For $z=4$, the solution can be written as
$q_{s}(J,K)|_{z=4}=\mathcal{S}(J,K)-(1/2)\sqrt{-4\mathcal{S}(J,K)^{2}+2\mathcal{H}(J,K)+\mathcal{Q}(J,K)/\mathcal{S}(J,K)},$ (C2)
where we have introduced the auxiliary functions
$\mathcal{H}(J,K)$ $\equiv\frac{3\cosh{(4J)}[\cosh{(4J)}+\cosh{(2K)}]}{\sinh^{2}{(2J)}[\cosh{(4J)}-2\sinh^{2}{(K)}]},$ (C3)
$\mathcal{Q}(J,K)$ $\equiv 16\left(\sum\nolimits_{n=\pm}[\tanh{(4J+nK)}-2\tanh{(2J+nK)}]\right)^{-1},$ (C4)
$\mathcal{S}(J,K)$ $\equiv(1/2)\sqrt{(2/3)\mathcal{H}(J,K)+(\mathcal{Q}(J,K)/6)\left(\mathcal{A}(J,K)+\Delta_{0}(J,K)/\mathcal{A}(J,K)\right)},$ (C5)
$\mathcal{A}(J,K)$ $\equiv 2^{-1/3}\left(\Delta_{1}(J,K)+\sqrt{\Delta^{2}_{1}(J,K)-4\Delta^{3}_{0}(J,K)}\right)^{1/3},$ (C6)
$\Delta_{0}(J,K)$ $\equiv-(3/4)\sum\nolimits_{n=\pm}[\tanh{(2J{+}nK)}+\tanh{(4J{+}nK)}]\sum\nolimits_{n=\pm}[\tanh{(2J{+}nK)}-\tanh{(4J{+}nK)}],$ (C7)
$\Delta_{1}(J,K)$ $\equiv\!216\cosh{(2J)}\!\cosh{(2K)}\!\operatorname{sech}{(4J{-}K)}\!\operatorname{sech}{(4J{+}K)}\!\sinh^{3}{(2J)}\sinh^{2}{(K)}[\cosh{(4J)}{+}\cosh{(2K)}]^{-2}.$ (C8)
Appendix D Oscillatory Instability
Here, we prove that $M_{2}(q_{s};J,K\neq 0)\neq 0$. To see this, we explicitly write out the first sum over $n$ in Eq. (18)
$M_{2}(q_{s};J,K)=\sum_{l=0}^{z}(\overline{\mathcal{P}}^{+}_{l}+\overline{\mathcal{P}}^{-}_{l})\left(\tanh{([2l{-}z]J{+}K)}{-}\tanh{([2l{-}z]J{-}K)}\right).$ (D1)
Note that $\overline{\mathcal{P}}^{\pm}_{l}(q_{s})>0$ for $q_{s}\in(-1,1)$, which follows straightforwardly from Eq. (19). Furthermore, since $\tanh{(x)}$ is an increasing function of $x$, we have $\tanh{([2l-z]J+K)}-\tanh{([2l-z]J-K)}>0$ for $K>0$ and $\tanh{([2l-z]J+K)}-\tanh{([2l-z]J-K)}<0$ for $K<0$. Hence, $M_{2}(q_{s};J,K)$ is given by a sum over strictly positive (for $K>0$) or negative (for $K<0$) terms, rendering $M_{2}(q_{s};J,K\neq 0)\neq 0$. This results in complex eigenvalues for $\lambda_{\pm}(q_{s};J,K)$ as shown in Fig. 1b.
Appendix E Mean-field approximation
@@
A less accurate technique to obtain approximate evolution equations is the mean-field (MF) approximation (originally developed in Penrose (1991)), where one makes the rudimentary (uncontrolled) assumption
$\langle\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\rangle\approx\tanh{\langle\Delta E^{\mu}_{i}/2\rangle},$ (E1)
@@
$\tau\frac{{\rm d}m^{a}(t)}{{\rm d}t}+m^{a}(t)$ $=\tanh{(zJ_{a}m^{a}(t)+K_{a}m^{b}(t))},$
$\tau\frac{{\rm d}m^{b}(t)}{{\rm d}t}+m^{b}(t)$ $=\tanh{(zJ_{b}m^{b}(t)+K_{b}m^{a}(t))},$ (E2)
which are exact on the fully connected mean-field lattice, where the local order is trivial (i.e. there is no sense of “local”), $q^{\mu\nu}(t)=m^{\mu}(t)m^{\nu}(t)$, and therefore $\mathcal{C}^{\mu\nu}(t)=0$. A linear stability analysis around the trivial steady state $m^{\mu}_{s}=0$ for $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$ leads to a linear stability equation where the eigenvalues of the linear stability matrix are given by
$\lambda^{\rm MF}_{\pm}(J,K)=(zJ-1)\pm iK.$ (E3)
Hence, the Hopf bifurcation occurs at $J=1/z$ and $K\neq 0$, such that ${\rm Re}(\lambda^{\rm MF}_{\pm})=0$ and ${\rm Im}(\lambda^{\rm MF}_{\pm})\neq 0$. This corresponds to a straight vertical line in the $(J,K)$-plane, as shown in Fig. E1a and also found in Avni et al. (2023). Notably, in the MF approximation we do not observe a critical value for local order, which is present in the more accurate BG approximation.
Figure E1: Phase diagram for global order $m^{\mu}_{s}$ obtained with the mean-field (MF) approximation (a) and the monomer approximation (b) for the perfect nonreciprocal setting with $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$. In the MF approximation the Hopf bifurcation (blue line) is set by $J=1/z$. In the monomer approximation, the Hopf bifurcations (blue line) is set by a critical local order $q_{s}=1/z$, which is more similar to the Bethe-Guggenheim approximation where $q_{s}=1/(z-1)$ (see Fig. 2d).
@@
MF (E2) $=m^{\mu}(t)m^{\nu}(t)$ $J{=}1/z$
monomer (8)+(F2) (9)-(10)+(F2); slaved by $m^{\mu}(t)$ $q_{s}{=}1/z$
BG (8)+(14) (9)-(10)+(14); not slaved $q_{s}{=}1/(z{-}1)$
Appendix F Monomer approximation
Another approximation technique we developed in this work is what we call the “monomer approximation”. It is more accurate than the MF but less accurate than the BG approximation.
The conceptual difference between the MF on the one hand, and the monomer and BG approximations on the other hand, lies in the treatment of the average $\langle\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\rangle$. Whereas the MF approximation simply moves the average to the argument as shown in Eq. (E1), the monomer and BG approximations use the fact that the value of $\Delta E^{\mu}_{i}/2$ lies in an enumerable set given by $U^{\mu}_{l,n}\equiv[2l-z]J_{\mu}+[2n-1]K_{\mu}$ with $l\in\{0,..,z\}$ and $n\in\{0,1\}$. This allows for an explicit summation
$\langle\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\rangle=\sum_{l=0}^{z}\sum_{n=0}^{1}\mathcal{P}^{\mu}_{l,n}(t)\tanh{(U^{\mu}_{l,n})},$ (F1)